# Scaling Gaussian Process Regression with Full Derivative Observations

### v1

@@ Abstract
We present a scalable Gaussian Process (GP) method that can fit and predict full derivative observations called DSoftKI. It extends SoftKI, a method that approximates a kernel via softmax interpolation from learned interpolation point locations, to the setting with derivatives. DSoftKI enhances SoftKI’s interpolation scheme to incorporate the directional orientation of interpolation points relative to the data. This enables the construction of a scalable approximate kernel, including its first and second-order derivatives, through interpolation. We evaluate DSoftKI on a synthetic function benchmark and high-dimensional molecular force field prediction (100-1000 dimensions), demonstrating that DSoftKI is accurate and can scale to larger datasets with full derivative observations than previously possible.
@@ 1 Introduction
To alleviate these challenges, scalable GP regression methods such as a Stochastic Variational GP (SVGP) (Hensman et al., 2013) have been extended to the setting with derivatives (DSVGP) (Padidar et al., 2021). It uses $m\ll n$ inducing points (Quinonero-Candela & Rasmussen, 2005; Snelson & Ghahramani, 2005) and achieves a time complexity of $\mathcal{O}(m^{3}d^{3})$ for posterior inference. Since the cubic scaling in $d$ can be prohibitive, a DSVGP with directional derivatives (DDSVGP) (Padidar et al., 2021) introduces $p\ll d$ inducing directions, an analogue of inducing points for dimensions, to achieve a time complexity of $\mathcal{O}(m^{3}p^{3})$ for posterior inference. [^1] This assumes $mp^{2}>d$ which no longer holds around $d=2000$ for $m=500$ and $p=2$. See Appendix C for details. However, this comes at the cost of directly predicting derivatives and introducing further approximations.
@@
In this paper, we present a GPwD method called DSoftKI [^2] Code can be found at https://github.com/danehuang/softki_gp_kit. that can fit and predict all derivative information (Section 4). It is an extension of SoftKI (Camaño & Huang, 2024), a scalable GP method that approximates a kernel via softmax interpolation from $m\ll n$ interpolation points whose locations are learned. Thus, it blends aspects of kernel interpolation (Wilson & Nickisch, 2015) which introduces kernel interpolation from a structured lattice with variational inducing point methods (Titsias, 2009) where inducing points are adapted to the dataset. To handle gradient information, DSoftKI modifies SoftKI’s interpolation scheme so that each interpolation point takes into account its directional orientation with respect to the underlying data. A scalable approximate kernel, including its first and second-order derivatives, can then be constructed via interpolation. This contrasts with methods such as DSVGP/DDSVP, which introduce separate inducing points for values and gradients as well as require computing kernel derivatives, leading to scaling challenges in both $n$ and $d$. As a result, DSoftKI achieves a time complexity of $\mathcal{O}(ndm^{2})$ for posterior inference. Because each datapoint consists of $d+1$ values to fit, we can view DSoftKI as achieving similar complexity to an approximate GP (e.g.(Titsias, 2009)) that is fit to $n(d+1)$ datapoints.
@@
We evaluate DSoftKI on synthetic functions with known derivatives as a baseline to compare against existing GP regression and GPwD regression methods. Since DSoftKI enables derivative modeling with GPs at larger $n$ and $d$ than previously possible, we also evaluate its efficacy on a high-dimensional molecular force field modeling task ($d$ around $100-1000$) which requires predicting gradients. Our experiments show that DSoftKI is a promising method for GPwD regression with full derivative prediction, both in terms of accuracy (e.g., see Figure 1) and scaling to larger $n$ and $d$ than previously possible (Section 5).
@@
Figure 1: Comparison of GPwD regression with vanilla GPwD, DSVGP, DDSVGP, and DSoftKI on the Branin surface (2D). We plot the contours of the surface, the gradient with respect to the first argument ($\nabla_{1}$), and the gradient with respect to the second argument ($\nabla_{2}$). Vanilla GPwD is accurate but intractable for sizable $n$ and/or $d$. DSVGP forms a nice approximation of the original surface but is intractable for large $d$. DDSVGP with uses $p=2$ directional derivatives to enhance scalability but loses fidelity in modeling the surface. DSoftKI provides an accurate and scalable approximation of the surface.
@@
Table 1: A comparison of GPwD methods across several dimensions. As a reminder, $n$ is the number of points, $d$ is the dimensionality of the data, $m$ is the number of inducing/interpolation points, and $p$ is the number of inducing directions. The proposed method, DSoftKI, is scalable, predicts full derivative observations, and does not require computing first or second-order derivatives of the kernel.
@@ 2 Related Work
One strategy for scaling GPwD regression is to extend an existing scalable GP method to the setting with derivatives. In this vein, both Stochastic Variational Gaussian Process (SVGP) (Hensman et al., 2013) and Structured Kernel Interpolation (SKI) (Wilson & Nickisch, 2015) have been extended to the setting with derivatives. Since we also follow this strategy, we briefly summarize these methods here and refer the reader to Appendix C for more background on these related works.
DSVGP (Padidar et al., 2021) extends SVGP to the setting with derivatives by modifying the SVGP variational approximation to the GPwD kernel, resulting in a kernel matrix of size $m(d+1)\times m(d+1)$ where $m$ is the number of inducing points (Snelson & Ghahramani, 2005; Quinonero-Candela & Rasmussen, 2005). Like SVGP, the inducing points in DSVGP are learned by optimizing an evidence lower bound (ELBO) that can be computed with stochastic variational inference. The time complexity of computing the ELBO for each minibatch of optimization has time complexity of $\mathcal{O}(m^{3}d^{3})$ and the time complexity of posterior inference is $\mathcal{O}(m^{3}d^{3})$. Since this can be prohibitive for large $d$, DDSVGP (Padidar et al., 2021) utilizes $p\ll d$ directional derivatives to further improve the time complexity of computing the ELBO per minibatch and posterior inference to $\mathcal{O}(m^{3}p^{3})$ since the $d$-dimensional gradients have been reduced to $p$-dimensional directional derivatives.
DSKI (Eriksson et al., 2018) extends SKI (Wilson & Nickisch, 2015) to the setting with derivatives by approximating the gradients of the SKI interpolation kernel. Like SKI, the resulting DSKI kernel has structure that enables fast matrix-vector multiplications (MVMs), and consequently, conjugate gradient (CG) methods to perform GP inference. Unlike SKI which uses a cubic interpolation scheme, DSKI uses a quintic interpolation scheme to better handle gradient observations, resulting in a time complexity of a single MVM of $O(nd6^{d}+m\log m)$. Thus, the scaling in the dimension $d$ is even worse than SKI, limiting the application of DSKI to even smaller dimensions. There are other kernel interpolation methods that improve the dimensionality scaling (e.g., see (Kapoor et al., 2021; Yadav et al., 2023)) that can be used besides SKI. However, to the best of our knowledge, these works have not been extended to the setting with derivatives.
@@ 3.1 Gaussian Processes
[…]
@@
Once the hyperparameters have been set, we can perform posterior inference. The posterior predictive distribution has the closed-form solution (Rasmussen & Williams, 2005)
@@ 3.2 Gaussian Processes with Derivative Information
If $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a GP with kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, then $\nabla f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is also a GP with kernel $k^{\prime}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{(d\times d)}$ defined as $k^{\prime}(x,y)=[\frac{\partial^{2}k}{\partial x_{i}\partial y_{j}}(x,y)]_{ij}=\nabla_{x}k(x,y)\nabla_{y}^{T}$. Consequently, we can construct a GP with derivative observations (GPwD), a random vector-valued function $\tilde{f}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d+1}$ defined as
@@
$\tilde{\mathbf{y}}\,|\,\tilde{f}(\mathbf{x})$ $\sim\mathcal{N}(\tilde{f}(\mathbf{x}),\tilde{\mathbf{\Lambda}})$ (likelihood)
where $\tilde{\mathbf{y}}=[y_{i}\,dy_{i}]_{i1}$ is a $n(d+1)\times 1$ vector of values and gradient labels and
@@
[…]
@@ 3.3 Soft Kernel Interpolation
Soft kernel interpolation (SoftKI) (Camaño & Huang, 2024) is an approximate GP method that uses an approximate kernel
@@
$\sigma_{\mathbf{z}}^{j}(x)=\frac{\exp\left(-\left\|x/T-z_{j}\right\|\right)}{\sum_{k=1}^{m}\exp\left(-\left\|x/T-z_{k}\right\|\right)}$ (9)
performs softmax interpolation between $m\ll n$ interpolation points $\mathbf{z}$ whose locations are learned, and $T\in\mathbb{R}^{d}$ is a learnable temperature vector akin to using automatic relevance detection (ARD) (MacKay et al., 1994) to set lengthscales for different dimensions. Thus, it is a method that combines aspects of kernel interpolation from SKI and adaptability of inducing point locations to data as in variational inducing points method to obtain a scalable GP method.
@@
$\log\bar{p}(\mathbf{y}\,|\,\mathbf{x};\theta)=-\frac{1}{2}\left[\mathbf{u}_{0}^{\top}\mathbf{D}_{\theta}\mathbf{u}_{0}+\frac{1}{l}\sum_{j=1}^{l}\mathbf{u}_{j}^{\top}(\mathbf{D}_{\theta}\mathbf{z}_{j})\right]$ (11)
where $\mathbf{D}_{\theta}=\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SoftKI}}(\ell,\gamma,T,\mathbf{z})+\mathbf{\Lambda}(\beta)$, $\mathbf{u}_{1},\dots,\mathbf{u}_{l}$ are solutions to the equations $\mathbf{D}_{\theta}(\mathbf{u}_{0}\,\mathbf{u}_{1}\dots\mathbf{u}_{l})=(\mathbf{y}\,\mathbf{z}_{1}\dots\mathbf{z}_{l})$, and $\mathbf{u}_{j}$ for $1\leq j\leq l$ are Gaussian random vectors normalized to have unit length. By construction,
$\nabla_{\theta}\log p(\mathbf{y}\,|\,\mathbf{x};\theta)\approx\nabla_{\theta}\log\tilde{p}(\mathbf{y}\,|\,\mathbf{x};\theta)$ (12)
so that the gradient of the Hutchinson’s pseudoloss is approximately equal to the gradient of the MLL. The time complexity of computing the Hutchinson’s pseudoloss per minibatch of size $b$ is $\mathcal{O}(b^{2}+m^{3})$.
@@
where $\hat{\mathbf{K}}_{\mathbf{x}\mathbf{z}}=\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}$ and $\mathbf{\hat{C}}=\mathbf{K_{\mathbf{z}\mathbf{z}}}+\mathbf{\hat{K}_{\mathbf{z}\mathbf{x}}}\mathbf{\Lambda}^{-1}\mathbf{\hat{K}_{\mathbf{x}\mathbf{z}}}$. It is similar to the posterior predictive of a Sparse Gaussian Process Regression (SGPR) (Titsias, 2009), the difference being that we replace kernels in the SGPR posterior with the interpolated versions. The time complexity of posterior inference is $\mathcal{O}(nm^{2})$.
@@ 4.1 Soft Kernel Interpolation with Derivatives
where we have abbreviated $\sigma^{j}_{\mathbf{z}}(x_{i})=\sigma_{ij}$ and $\nabla\sigma^{j}_{\mathbf{z}}(x_{i})=\partial\sigma_{ij}$ to reduce clutter. Notably, the computation of the DSoftKI kernel does not require computing the first-order or second-order derivatives of the kernel. Instead, it is approximated via interpolation. Consequently, it is tractable to perform hyperparameter optimization with first-order gradient methods. Additionally, it can support a wide range of kernels, including learned kernels (Wilson et al., 2016). This contrasts with methods such as DDSVGP where the kernel and its gradients are hard-coded in practice to alleviate the computational cost and to enable tractable hyperparameter optimization with gradient-based methods.
@@
While the DSoftKI kernel affords computational savings, it also increases the burden of kernel interpolation to approximate a kernel’s first and second-order derivatives. To solve this challenge, we adapt the softmax interpolation scheme to additionally account for directional orientation of interpolation points relative to the data. In this way, interpolation points that are close to each other can nevertheless model variations in surface curvature as influenced by relevant gradient observations. The original softmax interpolation scheme would be unable to distinguish these variations since the shared temperature parameter enforces a shared orientation for every interpolation point.
To implement the above idea, we associate each interpolation point $z_{k}$ with a corresponding learnable temperature vector $T_{k}\in\mathbb{R}^{d}$ as in
$\sigma_{\mathbf{z}}^{j}(x)=\frac{\exp\left(-\left\|x/T_{k}-z_{j}\right\|\right)}{\sum_{k=1}^{m}\exp\left(-\left\|x/T_{k}-z_{k}\right\|\right)}\,.$ (17)
This introduces $d\times m$ extra parameters $\mathbf{T}=(T_{1}\dots T_{m})^{(d\times m)}$ to the DSoftKI model that can be learned via hyperparameter optimization. The gradient of the softmax interpolation has closed-form solution
$\nabla_{x_{i}}\sigma_{\mathbf{z}}^{j}(x_{i})=\sum_{k=1}^{m}\sigma_{\mathbf{z}}^{j}(x_{i})(\delta_{jk}-\sigma_{\mathbf{z}}^{k}(x_{i}))\frac{x_{i}/T_{k}-z_{k}}{T_{k}\lVert x_{i}/T_{k}-z_{k}\rVert}$ (18)
which can be obtained in the standard way via the chain-rule. The term
$\frac{x_{i}/T_{k}-z_{k}}{T_{k}(\lVert x_{i}/T_{k}-z_{k}\rVert+\epsilon)}=\frac{1}{T_{k}^{2}}\frac{x_{i}-T_{k}z_{k}}{\lVert x_{i}-T_{k}z_{k}\rVert+\epsilon}\,$ (19)
where we have added a small factor $\epsilon>0$ to the denominator to avoid division by zero, is thus a unit vector with scale $1/T_{k}^{2}$ which gives the directional vector between each data point $x_{i}$ and point $T_{k}z_{k}$. As a result, points that are nearby can nevertheless have different influences on the interpolation strength of the gradients depending on their respective directions. Figure 3 illustrates the differences in surface reconstruction using a shared versus individual temperature vector across interpolation points.

### v2

@@ Abstract
We present a scalable Gaussian Process (GP) method called DSoftKI that can fit and predict full derivative observations. It extends SoftKI, a method that approximates a kernel via softmax interpolation, to the setting with derivatives. DSoftKI enhances SoftKI’s interpolation scheme by replacing its global temperature vector with local temperature vectors associated with each interpolation point. This modification allows the model to encode local directional sensitivity, enabling the construction of a scalable approximate kernel, including its first and second-order derivatives, through interpolation. Moreover, the interpolation scheme eliminates the need for kernel derivatives, facilitating extensions such as Deep Kernel Learning (DKL). We evaluate DSoftKI on synthetic benchmarks, a toy n-body physics simulation, standard regression datasets with synthetic gradients, and high-dimensional molecular force field prediction (100-1000 dimensions). Our results demonstrate that DSoftKI is accurate and scales to larger datasets with full derivative observations than previously possible.
@@ 1 Introduction
To alleviate these challenges, scalable GP regression methods such as a Stochastic Variational GP (SVGP) (Hensman et al., 2013) have been extended to the setting with derivatives (DSVGP) (Padidar et al., 2021). It uses $m\ll n$ inducing points (Quinonero-Candela and Rasmussen, 2005; Snelson and Ghahramani, 2005) and achieves a time complexity of $\mathcal{O}(m^{3}d^{3})$ for posterior inference. Since the cubic scaling in $d$ can be prohibitive, a DSVGP with directional derivatives (DDSVGP) (Padidar et al., 2021) introduces $p\ll d$ inducing directions, an analogue of inducing points for dimensions, to achieve a time complexity of $\mathcal{O}(m^{3}p^{3})$ for posterior inference. [^1] This assumes $mp^{2}>d$ which no longer holds around $d=2000$ for $m=500$ and $p=2$. See Appendix E for details. However, this comes at the cost of directly predicting derivatives and introducing further approximations.
@@
In this paper, we present a GPwD method called DSoftKI [^2] Code available at https://github.com/base26labs/dsoftki_gp. that can fit and predict all derivative information (Section 4). It is an extension of SoftKI (Camaño and Huang, 2025), a scalable GP method that approximates a kernel via softmax interpolation from $m\ll n$ interpolation points whose locations are learned. Thus, it blends aspects of kernel interpolation (Wilson and Nickisch, 2015) which introduces kernel interpolation from a structured lattice with variational inducing point methods (Titsias, 2009) where inducing points are adapted to the dataset. To handle gradient information, DSoftKI modifies SoftKI’s interpolation scheme by replacing its global temperature vector with local temperature vectors associated with each interpolation point. This allows the model to encode local directional sensitivity (Section 4.1), enabling the construction of a scalable approximate kernel, including its first and second-order derivatives, through interpolation. Notably, this design eliminates the need for kernel derivatives, facilitating extensions such as Deep Kernel Learning (DKL) (Wilson et al., 2016) which use learned kernels (see Appendix D). This contrasts with methods such as DSVGP/DDSVP, which introduce separate inducing points for values and gradients as well as require computing kernel derivatives, leading to scaling challenges in both $n$ and $d$ (see Appendix E for more details). As a result, DSoftKI achieves a time complexity of $\mathcal{O}(m^{2}nd)$ for posterior inference. Because each datapoint consists of $d+1$ values to fit, we can view DSoftKI as achieving similar complexity to an approximate GP (e.g.(Titsias, 2009)) that is fit to $n(d+1)$ datapoints.
@@
We evaluate DSoftKI on synthetic functions with known derivatives as a baseline to compare against existing GP regression and GPwD regression methods. Since DSoftKI enables derivative modeling with GPs at larger $n$ and $d$ than previously possible, we also evaluate its efficacy on a high-dimensional molecular force field modeling task ($d$ around 1001000100-1000) which requires predicting gradients. To test a broader range of datasets, we also validate performance on a toy n-body physics simulation and standard UCI regression benchmarks with synthetic derivatives (Appendix C). Our experiments show that DSoftKI is a promising method for GPwD regression with full derivative prediction, both in terms of accuracy (e.g., see Figure 1) and scaling to larger $n$ and $d$ than previously possible (Section 5).
@@
Figure 1: Comparison of GPwD regression with vanilla GPwD, DSVGP, DDSVGP, and DSoftKI on the Branin surface (2D). We plot the contours of the surface, the gradient with respect to the first argument ($\nabla_{1}$), and the gradient with respect to the second argument ($\nabla_{2}$). Vanilla GPwD is accurate but intractable for sizable $n$ and/or $d$. DSVGP forms a nice approximation of the original surface but is intractable for large $d$. DDSVGP uses $p=2$ directional derivatives to enhance scalability but loses fidelity in modeling the surface. DSoftKI provides an accurate and scalable approximation of the surface.
@@
Table 1: A comparison of GPwD methods across several dimensions. As a reminder, $n$ is the number of points, $d$ is the dimensionality of the data, $m$ is the number of inducing/interpolation points, and $p$ is the number of inducing directions. The proposed method, DSoftKI, is scalable, predicts full derivative observations, and does not require computing first or second-order derivatives of the kernel to construct the GPwD kernel. The kernel approximations made by each method are discussed in more detail in Appendix A.1.
@@ 2 Related Work
One strategy for scaling GPwD regression is to extend an existing scalable GP method to the setting with derivatives. In this vein, both Stochastic Variational Gaussian Process (SVGP) (Hensman et al., 2013) and Structured Kernel Interpolation (SKI) (Wilson and Nickisch, 2015) have been extended to the setting with derivatives. Since we also follow this strategy, we briefly summarize these methods here and refer the reader to Appendix E for more background on these related works.
DSVGP (Padidar et al., 2021) extends SVGP to the setting with derivatives by modifying the SVGP variational approximation to the GPwD kernel, resulting in a kernel matrix of size $m(d+1)\times m(d+1)$ where $m$ is the number of inducing points (Snelson and Ghahramani, 2005; Quinonero-Candela and Rasmussen, 2005). Like SVGP, the inducing points in DSVGP are learned by optimizing an evidence lower bound (ELBO) that can be computed with stochastic variational inference. The time complexity of computing the ELBO for each minibatch of optimization has time complexity of $\mathcal{O}(m^{3}d^{3})$ and the time complexity of posterior inference is $\mathcal{O}(m^{3}d^{3})$. Since this can be prohibitive for large $d$, DDSVGP (Padidar et al., 2021) utilizes $p\ll d$ directional derivatives to further improve the time complexity of computing the ELBO per minibatch and posterior inference to $\mathcal{O}(m^{3}p^{3})$ since the $d$-dimensional gradients have been reduced to $p$-dimensional directional derivatives.
DSKI (Eriksson et al., 2018) extends SKI (Wilson and Nickisch, 2015) to the setting with derivatives by approximating the gradients of the SKI interpolation kernel. Like SKI, the resulting DSKI kernel has structure that enables fast matrix-vector multiplications (MVMs), and consequently, conjugate gradient (CG) methods to perform GP inference. Unlike SKI which uses a cubic interpolation scheme, DSKI uses a quintic interpolation scheme to better handle gradient observations, resulting in a time complexity of a single MVM of $O(nd6^{d}+m\log m)$. Thus, the scaling in the dimension $d$ is even worse than SKI, limiting the application of DSKI to even smaller dimensions. There are other kernel interpolation methods that improve the dimensionality scaling (e.g., see Kapoor et al. (2021); Yadav et al. (2023)) that can be used besides SKI. However, to the best of our knowledge, these works have not been extended to the setting with derivatives.
@@ 3.1 Gaussian Processes
$\log p(\mathbf{y}\,|\,\mathbf{x};\theta)=\mathcal{N}(\mathbf{y}\,|\,\mathbf{0},\mathbf{K}_{\mathbf{x}\mathbf{x}}(\ell,\gamma)+\mathbf{\Lambda}(\beta))$ (1)
@@
Once the hyperparameters have been set, we can perform posterior inference. The posterior predictive distribution has the closed-form solution (Rasmussen and Williams, 2005)
@@ 3.2 Gaussian Processes with Derivative Information
If $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a GP with kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, then $\nabla f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is also a GP with kernel $k^{\prime}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{(d\times d)}$ defined as
$k^{\prime}(x,y)=\left[\frac{\partial^{2}k(x,y)}{\partial x_{i}\partial y_{j}}\right]_{ij}=\nabla_{x}k(x,y)\nabla_{y}^{T}\,.$ (3)
Consequently, we can construct a GP with derivative observations (GPwD), a random vector-valued function $\tilde{f}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d+1}$ defined as
@@
[…]
where $\tilde{\mathbf{y}}=[y_{i}\,dy_{i}^{\top}]_{1i}^{\top}$ is a $n(d+1)\times 1$ vector of values and gradient labels and
@@
$\log p(\mathbf{y}\,|\,\mathbf{x};\theta)$ $=\mathcal{N}(\mathbf{y}\,|\,0,\tilde{\mathbf{K}}_{\mathbf{x}\mathbf{x}}(\ell,\gamma)+\tilde{\mathbf{\Lambda}}(\beta_{v},\beta_{g}))$ (7)
@@ 3.3 Soft Kernel Interpolation
Soft kernel interpolation (SoftKI) (Camaño and Huang, 2025) is an approximate GP method that uses an approximate kernel
@@
$\sigma_{\mathbf{z}}^{j}(x)=\frac{\exp\left(-\left\|x\oslash T-z_{j}\right\|\right)}{\sum_{k=1}^{m}\exp\left(-\left\|x\oslash T-z_{k}\right\|\right)}$ (10)
[…]
@@
$\log\bar{p}(\mathbf{y}\,|\,\mathbf{x};\theta)=-\frac{1}{2}\left[\mathbf{u}_{0}^{\top}\mathbf{D}_{\theta}\mathbf{u}_{0}-\frac{1}{l}\sum_{j=1}^{l}\mathbf{u}_{j}^{\top}(\mathbf{D}_{\theta}\mathbf{w}_{j})\right]$ (12)
where $\mathbf{D}_{\theta}=\mathbf{K}_{\mathbf{x}\mathbf{x}}^{\text{SoftKI}}(\ell,\gamma,T,\mathbf{z})+\mathbf{\Lambda}(\beta)$, $\mathbf{u}_{1},\dots,\mathbf{u}_{l}$ are solutions to the equations $\mathbf{D}_{\theta}(\mathbf{u}_{0}\,\mathbf{u}_{1}\dots\mathbf{u}_{l})=(\mathbf{y}\,\mathbf{w}_{1}\dots\mathbf{w}_{l})$, and $\mathbf{w}_{j}$ for $1\leq j\leq l$ are Gaussian random vectors normalized to have unit length. Observe that
$\nabla_{\theta}\log p(\mathbf{y}\,|\,\mathbf{x};\theta)$ $=-\frac{1}{2}\Big[-\mathbf{u}_{0}^{\top}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\mathbf{u}_{0}+\operatorname{tr}\Big(\mathbf{D}_{\theta}^{-1}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\Big)\Big]$ (13)
$\approx-\frac{1}{2}\Big[-\mathbf{u}_{0}^{\top}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\mathbf{u}_{0}+\frac{1}{l}\sum_{j=1}^{l}\mathbf{u}_{j}^{\top}\frac{\partial\mathbf{D}_{\theta}}{\partial\mathbf{\theta}}\mathbf{w}_{j}\Big]$ (14)
$=\nabla_{\theta}\log\tilde{p}(\mathbf{y}\,|\,\mathbf{x};\theta)$ (15)
so that the gradient of the Hutchinson’s pseudoloss is approximately equal to the gradient of the MLL when approximated with Hutchinson’s stochastic trace estimator (Girard, 1989; Hutchinson, 1989). The time complexity of computing the Hutchinson’s pseudoloss per minibatch of size $b$ is $\mathcal{O}(b^{2}+m^{3})$.
@@
where $\hat{\mathbf{K}}_{\mathbf{x}\mathbf{z}}=\mathbf{\Sigma}_{\mathbf{x}\mathbf{z}}\mathbf{K}_{\mathbf{z}\mathbf{z}}$ and $\mathbf{\hat{C}}=\mathbf{K_{\mathbf{z}\mathbf{z}}}+\mathbf{\hat{K}_{\mathbf{z}\mathbf{x}}}\mathbf{\Lambda}^{-1}\mathbf{\hat{K}_{\mathbf{x}\mathbf{z}}}$. It is similar to the posterior predictive of a Sparse Gaussian Process Regression (SGPR) (Titsias, 2009), the difference being that we replace kernels in the SGPR posterior with the interpolated versions. The time complexity of posterior inference is $\mathcal{O}(m^{2}n)$.
@@ 4.1 Soft Kernel Interpolation with Derivatives
where we have abbreviated $\sigma^{j}_{\mathbf{z}}(x_{i})=\sigma_{ij}$ and $\nabla\sigma^{j}_{\mathbf{z}}(x_{i})=\partial\sigma_{ij}$ to reduce clutter. Notably, the computation of the DSoftKI kernel does not require computing the first-order or second-order derivatives of the kernel. Instead, it is approximated via interpolation. Consequently, it is tractable to perform hyperparameter optimization with first-order gradient methods. Additionally, it can support a wide range of kernels, including learned kernels as in DKL. This contrasts with methods such as DDSVGP where the kernel and its gradients are hard-coded in practice to alleviate the computational cost and to enable tractable hyperparameter optimization with gradient-based methods. We refer the reader to Appendix D for a demonstration of DKL with DSoftKI.
@@
While the DSoftKI kernel affords computational savings, it also increases the burden of kernel interpolation to approximate a kernel’s first and second-order derivatives. To solve this challenge, we adapt the softmax interpolation scheme to additionally account for directional orientation of interpolation points relative to the data. More concretely, we replace the global temperature vector with local temperature vectors associated with each interpolation point. In this way, interpolation points that are close to each other can nevertheless model variations in surface curvature as influenced by relevant gradient observations. The original softmax interpolation scheme would be unable to distinguish these variations since the global temperature vector enforces a shared orientation for every interpolation point.
To implement the above idea, we associate each interpolation point $z_{k}$ with a corresponding learnable temperature vector $T_{k}\in\mathbb{R}^{d}$ for $1\leq k\leq m$ as in
$\sigma_{\mathbf{z}}^{j}(x)=\frac{\exp\left(-\left\|x\oslash T_{j}-z_{j}\right\|\right)}{\sum_{k=1}^{m}\exp\left(-\left\|x\oslash T_{k}-z_{k}\right\|\right)}\,.$ (20)
This introduces $d\times m$ extra parameters $\mathbf{T}^{(d\times m)}=(T_{1}\dots T_{m})$ to the DSoftKI model that can be learned via hyperparameter optimization. The gradient of the softmax interpolation has closed-form solution
$\nabla_{x_{i}}\sigma_{\mathbf{z}}^{j}(x_{i})=-\sum_{k=1}^{m}\sigma_{\mathbf{z}}^{j}(x_{i})(\delta_{jk}-\sigma_{\mathbf{z}}^{k}(x_{i}))\frac{x_{i}\oslash T_{k}-z_{k}}{\lVert x_{i}\oslash T_{k}-z_{k}\rVert}\oslash T_{k}$ (21)
which can be obtained in the standard way via the chain-rule. The $j$-th component of the term
$\left[\frac{x_{i}\oslash T_{k}-z_{k}}{(\lVert x_{i}\oslash T_{k}-z_{k}\rVert+\epsilon)}\oslash T_{k}\right]_{j}=\frac{x_{ij}-T_{kj}z_{kj}}{T_{kj}^{2}(\lVert x_{i}\oslash T_{k}-z_{k}\rVert+\epsilon)}$ (22)
where we have added a small factor $\epsilon>0$ to the denominator to avoid division by zero, is proportional to the direction from the scaled interpolation point $T_{k}\odot z_{k}$ to the data point $x_{i}$ along dimension $j$, with scaling factor $1/T_{kj}^{2}$. As a result, points that are nearby can nevertheless have different influences on the interpolation strength of the gradients depending on their respective directions. Figure 3 illustrates the differences in surface reconstruction using a shared versus individual temperature vector across interpolation points. We refer the reader to Appendix B.5 for further discussion on the role of temperatures and their connection with lengthscales.